OpenStudy (idkwut):

The velocity of a particle on the x-axis is given by the differential equation. dx/dt = 3/4t^(1/2). The particle is at x = 5, when t = 1. The position of the particle as a function of time is x(t) = ___.

OpenStudy (anonymous):

Do you mean \( \frac{dx} {dt}\)

OpenStudy (idkwut):

Oh sorry, yea i do

OpenStudy (anonymous):

\[ x(t)=\int \frac{3 \sqrt{t}}{4} \, dt=\frac{t^{3/2}}{2}+C\\ 5=x(1)= \frac 1 2 +C\\ C=\frac 92\\ x(t)=\int \frac{3 \sqrt{t}}{4} \, dt=\frac{t^{3/2}}{2}+\frac 9 2\\ \]